# Preference-based Reinforcement Learning beyond Pairwise Comparisons: Benefits of Multiple Options

### v1

@@ 3.1 Loss Functions and Rank Breaking
Here, $\ell^{(j)}_{t}(\bm{\theta})$ denotes the negative log-likelihood loss under the Multinomial Logit (MNL) model [48], conditioned on the assortment being the remaining actions in $S_{t}$ after removing the previously selected actions $\sigma_{t1},\dots,\sigma_{t(j-1)}$}—that is, over the set $S_{t}\setminus\{\sigma_{t1},\dots,\sigma_{t(j-1)}\}$.
Rank-Breaking (RB) loss. In addition to this standard approach, one can replace the full $|S_{t}|$-action ranking with its $\binom{|S_{t}|}{2}$ pairwise comparisons. This technique, referred to as rank breaking (RB), decomposes (partial) ranking data into individual pairwise comparisons, treating each comparison as independent [6, 34, 32, 69]. Thus, the RB loss is defined as:
@@
This approach is applied in the current RLHF for LLM (e.g., Ouyang et al. [59]) and is also studied in the theoretical RLHF paper [96] under the offline setting.
@@ 3.2 Online Parameter Estimation
Motivated by recent advances in Multinomial Logit (MNL) bandits [95, 41, 43], we adopt an online mirror descent (OMD) algorithm to estimate the underlying parameter $\bm{\theta}^{\star}$, instead of relying on maximum likelihood estimation (MLE). This enables a constant per-round computational cost, in contrast to the MLE-based approach, whose cost grows linearly with the number of rounds $t$.
@@
The optimization problem (5) can be solved using a single projected gradient step [57], which enjoys a computational cost of only $\mathcal{O}(Kd^{3})$—independent of $t$ [50], unlike MLE—and requires only $\mathcal{O}(d^{2})$ storage, thanks to the incremental updates of $\tilde{H}_{t}^{(j)}$ and $H_{t}$.
@@
where we set $\widehat{\bm{\theta}}_{t}^{(j,|S_{t}|+1)}=\widehat{\bm{\theta}}_{t}^{(j+1,j+2)}$ for all $j<|S_{t}|-1$ and for the final pair, let $\widehat{\bm{\theta}}_{t}^{(|S_{t}|-1,|S_{t}|+1)}=\widehat{\bm{\theta}}_{t+1}^{(1,2)}$. Also, the matrix $\tilde{H}_{t}^{(j,k)}$ is defined as $\tilde{H}_{t}^{(j,k)}\!:=H_{t}+\eta\sum_{(j^{\prime},k^{\prime})\leq(j,k)}\!\ \nabla^{2}\ell_{t}^{(j^{\prime},k^{\prime})}(\widehat{\bm{\theta}}^{(j^{\prime},k^{\prime})}_{t})$  [^2] We write $(j^{\prime},k^{\prime})\leq(j,k)$ to indicate lexicographic order, i.e., $j^{\prime}<j$ or $j^{\prime}=j$ and $k^{\prime}\leq k$. , where
@@ 4 M-AUPO: Maximizing Average Uncertainty
In this section, we propose a new algorithm, M-AUPO, designed to select an assortment that maximizes average uncertainty of $S_{t}$, thereby leveraging the potential benefits of a large $K$. At each round $t$, a context $x_{t}$ is drawn from a fixed but unknown distribution $\rho$. The algorithm then selects a reference action-assortment pair $(\bar{a}_{t},S_{t})$ by maximizing the average feature uncertainty—measured in the $H_{t}^{-1}$-norm—relative to a candidate reference action $\bar{a}$ (Line 4):
$(\bar{a}_{t},S_{t})=\mathop{\mathrm{argmax}}_{\bar{a}\in\mathcal{A}}\mathop{\mathrm{argmax}}_{\begin{subarray}{c}S\in{\mathcal{S}}\\ \bar{a}\in S\end{subarray}}\frac{1}{|S|}\sum_{a\in S}\left\|\phi\left(x_{t},a\right)-\phi(x_{t},\bar{a})\right\|_{H_{t}^{-1}}.$ (9)
By construction, the reference action $\bar{a}_{t}$ is always included in the selected assortment $S_{t}$. This selection strategy plays a key role in our algorithm, as it promotes rapid reduction in reward uncertainty—particularly when the assortment size $|S_{t}|$ is large—by encouraging informative comparisons centered around the reference action. Importantly, the assortment selection rule in Equation (9) can be computed efficiently, without enumerating all $\binom{N}{K}$ possible subsets.
Remark 2 (Computational cost of $S_{t}$-selection).
The optimization in Equation (9) can be efficiently solved with a computational cost of $\tilde{\mathcal{O}}(N^{2}K)$ (see Appendix B for details). Furthermore, in Appendix H.1, we will show that the reference action $\bar{a}_{t}$ can be chosen arbitrarily, which further reduces the computational cost to $\tilde{\mathcal{O}}(NK)$.
Then, we observe the ranking feedback $\sigma_{t}$ from a labeler and update the parameter according to Procedure 1 if using the PL loss, or Procedure 2 if using the RB loss (Line 6). After $T$ rounds, the algorithm returns the final policy $\widehat{\pi}_{T}$, which selects actions by maximizing the estimated reward under the final parameter estimate, i.e., $\widehat{\pi}_{T}(x):=\mathop{\mathrm{argmax}}_{a}\phi(x,a)^{\top}\widehat{\bm{\theta}}_{T+1}$ (Line 7).
@@ 5 Main Results
In this section, we present our main theoretical contributions. In Section 5.1, we show that M-AUPO achieves a suboptimality gap that decreases with the size of the presented assortment $|S_{t}|$, implying improved performance when larger action subsets are offered for ranking feedback. In Section 5.2, we establish the near-matching lower bound.
@@ 5.1 Suboptimality Gap of M-AUPO
We begin by presenting the online confidence bound for the PL loss, derived by extending the results of Lee and Oh [43], who analyzed the MNL model [48]. Since the PL model constructs ranking probabilities as a product of MNL probabilities, their confidence bound can be directly applied to our setting by replacing the round $t$ with the cumulative number of updates $\sum_{s=1}^{t}|S_{s}|$.
@@
This confidence bound is free of any polynomial dependency on $K$, which is primarily made possible by the improved self-concordant-like properties proposed by Lee and Oh [43]. Moreover, for the RB loss, we can derive a confidence bound of the same order (see Corollary E.1). Based on this confidence bound, we derive the suboptimality gap for M-AUPO, with the proof deferred to Appendix D.
Theorem 1.
Let $\delta\in(0,1]$. We set $\lambda=\Omega\big(d\log(KT/\delta)+\eta(B+d)\big)$ and $\eta=\frac{1}{2}(1+3\sqrt{2}B)$. Define $\kappa:=e^{-4B}$. If Assumption 1 holds, then, with probability at least $1-\delta$, M-AUPO (Algorithm 3) achieves the following suboptimality gap:
$\textup{{SubOpt}}(T)=\tilde{\mathcal{O}}\left(\frac{d}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}+\frac{d^{2}K^{2}}{\kappa T}\right).$
Discussion of Theorem 1. For sufficiently large $T$, the second (non-leading) term becomes negligible, and Theorem 1 shows that the suboptimality gap of M-AUPO decreases as the assortment size $|S_{t}|$ increases. This establishes a strict advantage of receiving ranking feedback over larger assortments. Moreover, our result does not involve any $\mathcal{O}(e^{B})$ dependency in the leading term, a harmful dependency that commonly appears in prior works [68, 72, 96, 89, 94, 19, 79, 39]. Although very recent studies [20, 14] also achieve $\mathcal{O}(e^{B})$-free performance in the leading term, they rely on auxiliary techniques and are restricted to pairwise preference feedback. To the best of our knowledge, this is the first theoretical study that simultaneously establishes (i) the performance benefits of utilizing richer ranking feedback over larger assortments, and (ii) the elimination of the $\mathcal{O}(e^{B})$ dependence in the leading term of the PbRL framework when accommodating multiple (i.e., more than two) options.
Proof sketch of Theorem 1. We provide a proof sketch of Theorem 1. For simplicity, the main paper assumes that the term $\big\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\big\|_{H_{t}^{-1}}$ remains sufficiently small for all $a$ during all rounds. This is justified because the regret incurred in the rounds where this condition fails is bounded by lower-order terms and thus has a negligible impact (see Lemma D.4 and D.5).
@@
$\textup{{SubOpt}}(T)$ $=\frac{1}{T}\sum_{t=1}^{T}\underbrace{\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\bm{\theta}^{\star}}_{\text{realized regret of $\widehat{\pi}_{T}$ at round }t}+\underbrace{\frac{1}{T}\sum_{t=1}^{T}\text{MDS}_{t}}_{=\tilde{\mathcal{O}}(1/\sqrt{T})}$
@@
$\lesssim\frac{1}{T}\sum_{t=1}^{T}\big\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big\|_{H_{t}^{-1}}\underbrace{\big\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\big\|_{H_{T}}}_{=\tilde{\mathcal{O}}(\sqrt{d})}+\tilde{\mathcal{O}}\left(\frac{1}{\sqrt{T}}\right).$
@@
In the inequality, we first use the fact that $\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)^{\top}\big(\widehat{\bm{\theta}}_{T+1}-\bm{\theta}^{\star}\big)\geq 0$, which follows from definition of $\widehat{\pi}_{T}$, and then apply Hölder’s inequality together with the inequality $H_{T+1}\succeq H_{t}$. We now apply Corollary 1 to upper bound $\big\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\big\|_{H_{T}}$ by $\tilde{\mathcal{O}}(\sqrt{d})$. Next, using our assortment selection rule (9), we can bound $\big\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big\|_{H_{t}^{-1}}$ as follows:
$\frac{1}{2}\,\big\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big\|_{H_{t}^{-1}}\leq\frac{1}{|S_{t}|}\sum_{a\in S_{t}}\big\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\big\|_{H_{t}^{-1}}.$ (10)
Hence, the performance is expected to improve as the subset size $|S_{t}|$ increases.
2) Avoiding $\mathcal{O}(e^{B})$ by matrix concentration. To further bound the right-hand side of Equation (10), we first express $H_{t}$ as follows:
$H_{t}=\frac{1}{2}\sum_{s=1}^{t-1}\sum_{j=1}^{|S_{s}|}\mathbb{E}_{(a,a^{\prime})\sim P_{s}^{(j)}\times P_{s}^{(j)}}\left[\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)^{\top}\right]+\lambda\mathbf{I}_{d},$
where $P_{s}^{(j)}$ denote the (true) MNL distribution over the remaining actions in $S_{s}$ after removing the first j1j-1 selected actions, i.e., $P_{s}^{(j)}(a):=\frac{\exp\left(\phi(x_{s},a)^{\top}\bm{\theta}^{\star}\right)}{\sum_{a^{\prime}\in S^{(j)}_{s}}\exp\left(\phi(x_{s},a^{\prime})^{\top}\bm{\theta}^{\star}\right)}$, where $a\in S^{(j)}_{s}:=\{\sigma_{sj},\dots,\sigma_{s|S_{s}|}\}$. Furthermore, we define the regularized sample covariance matrix of feature differences, $\Lambda_{t}$, which, unlike $H_{t}$, does not incorporate local information:
$\Lambda_{t}:=\sum_{s=1}^{t-1}\sum_{a\in S_{s}}\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}.$
Aside from the very recent works [20, 14], which avoid the $\mathcal{O}(e^{B})$ dependency but only in pairwise comparison settings, most previous works on linear contextual dueling bandits and PbRL [68, 72, 96, 89, 94, 19, 79, 39] exhibit performance (either in terms of cumulative regret or suboptimality gap) that depends on $\mathcal{O}(Ke^{B})$ (or $\mathcal{O}(e^{B})$ in the case of pairwise comparisons). This dependency arises because these works apply a crude lower bound on $H_{t}$ by using the inequality $P_{s}^{(j)}(a)P_{s}^{(j)}(a^{\prime})\gtrsim\frac{1}{K^{2}e^{2B}}$. As a result, they derive $H_{t}\succsim\frac{1}{K^{2}e^{2B}}\Lambda_{t}$, which further implies $\big\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\big\|_{H_{t}^{-1}}\lesssim Ke^{B}\big\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\big\|_{\Lambda_{t}^{-1}}$. This leads directly to performance bounds that scale with $\mathcal{O}(Ke^{B})$.
To tackle this problem, we leverage the concentration lemma for covariance matrices (Corollary F.1) and for PSD matrices (Lemma F.4). The following lemma shows that, even without introducing additional algorithmic complexities specifically designed to avoid the $\mathcal{O}(e^{B})$ dependency—as done in Di et al. [20] and Chen et al. [14]—the standard analysis techniques are sufficient to eliminate the $\mathcal{O}(Ke^{B})$ dependency. In particular, we establish that $H_{t}$ approximates $\Lambda_{t}$ up to a constant factor.
Lemma 1.
Let $\lambda=\Omega(d\log(KT/\delta))$. Then, with probability at least $1-\delta$, we have
$H_{t}\succeq\frac{1}{50}\Lambda_{t}.$
Remark 3 (Applicability of Lemma 1).
Lemma 1 is expected to readily apply to most existing PbRL (or RLHF) and dueling bandit algorithms without requiring any modification to their original formulations, thereby eliminating the $\mathcal{O}(e^{B})$ dependency in the leading term.
By applying Lemma 1 and Cauchy-Schwartz inequality, we obtain:
$\frac{1}{T}\sum_{t=1}^{T}\frac{1}{|S_{t}|}\sum_{a\in S_{t}}\big\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\big\|_{H_{t}^{-1}}$ $\lesssim\frac{1}{T}\sqrt{\sum_{t=1}^{T}\frac{|S_{t}|}{|S_{t}|^{2}}}\underbrace{\sqrt{\sum_{a\in S_{t}}\big\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\big\|_{\Lambda_{t}^{-1}}^{2}}}_{=\tilde{\mathcal{O}}(\sqrt{d})}.$
Finally, applying the elliptical potential lemma (Lemma D.3), we concludes the proof.
Furthermore, we establish a similar suboptimality gap when using the RB loss (4) in place of the PL loss (3). The proof is provided in Appendix E.
Theorem 2.

### v2

@@ 3.1 Loss Functions and Rank-Breaking
Here, $\ell^{(j)}_{t}(\bm{\theta})$ denotes the negative log-likelihood loss under the Multinomial Logit (MNL) model [mcfadden1977modelling], conditioned on the assortment being the remaining actions in $S_{t}$ after removing the previously selected actions $\sigma_{t1},\dots,\sigma_{t(j-1)}$}—that is, over the set $S_{t}\setminus\{\sigma_{t1},\dots,\sigma_{t(j-1)}\}$.
Rank-Breaking (RB) loss. In addition to this standard approach, one can replace the full $|S_{t}|$-action ranking with its $\binom{|S_{t}|}{2}$ pairwise comparisons. This technique, referred to as rank-breaking (RB), decomposes (partial) ranking data into individual pairwise comparisons, treating each comparison as independent [azari2013generalized, khetan2016data, jang2017optimal, saha2024stop]. Thus, the RB loss is defined as:
@@
This approach is applied in the current RLHF for LLM (e.g., ouyang2022training) and is also studied in the theoretical RLHF paper [zhu2023principled] under the offline setting.
@@ 3.2 Online Parameter Estimation
Motivated by recent advances in Multinomial Logit (MNL) bandits [zhang2024online, lee2024nearly, lee2025improved], we adopt an online mirror descent (OMD) algorithm to estimate the underlying parameter $\bm{\theta}^{\star}$, instead of relying on maximum likelihood estimation (MLE). This enables a constant per-round computational cost, in contrast to the MLE-based approach, whose cost grows linearly with the number of rounds $t$.
@@
The optimization problem (5) can be solved using a single projected gradient step [orabona2019modern], which enjoys a computational cost of only $\mathcal{O}(Kd^{3})$—independent of $t$ [mhammedi2019lipschitz], unlike MLE—and requires only $\mathcal{O}(d^{2})$ storage, thanks to the incremental updates of $\tilde{H}_{t}^{(j)}$ and $H_{t}$.
@@
where we set $\widehat{\bm{\theta}}_{t}^{(j,|S_{t}|+1)}=\widehat{\bm{\theta}}_{t}^{(j+1,j+2)}$ for all $j<|S_{t}|-1$ and for the final pair, let $\widehat{\bm{\theta}}_{t}^{(|S_{t}|-1,|S_{t}|+1)}=\widehat{\bm{\theta}}_{t+1}^{(1,2)}$. The matrix $\tilde{H}_{t}^{(j,k)}$ is defined as $\tilde{H}_{t}^{(j,k)}\!:=H_{t}+\eta\sum_{(j^{\prime},k^{\prime})\leq(j,k)}\!\ \nabla^{2}\ell_{t}^{(j^{\prime},k^{\prime})}(\widehat{\bm{\theta}}^{(j^{\prime},k^{\prime})}_{t})$  [^2] We write $(j^{\prime},k^{\prime})\leq(j,k)$ to indicate lexicographic order, i.e., $j^{\prime}<j$ or $j^{\prime}=j$ and $k^{\prime}\leq k$. , where
@@ 4 M-AUPO: Maximizing Average Uncertainty
In this section, we propose a new algorithm, M-AUPO, which selects an assortment that maximizes the average uncertainty of $S_{t}$, thereby exploiting the potential benefits of a larger $K$. For clarity of presentation, we first define the MNL probability [mcfadden1977modelling] for a given assortment $S$ at round $t$ as follows:
$P_{t}(a|S;\bm{\theta}):=\frac{\exp\left(\phi(x_{t},a)^{\top}\bm{\theta}\right)}{\sum_{a^{\prime}\in S}\exp\left(\phi(x_{t},a^{\prime})^{\top}\bm{\theta}\right)},\quad\forall a\in S.$ (9)
Given an assortment $S$ and ranking feedback $\sigma$, let $S^{(j)}_{\sigma}:=\{\sigma_{j},\dots,\sigma_{|S|}\}$ denote the remaining actions in $S$ after removing the first j1j-1 actions. Let $\mathbb{P}_{t}(\sigma|S;\bm{\theta})=\mathbb{P}(x_{t},\sigma|S;\bm{\theta})$, for simplicity. Then, the PL model in Equation (1) can be expressed as follows:
$\mathbb{P}_{t}(\sigma|S;\bm{\theta})$ $=P_{t}(\sigma_{1}|S;\bm{\theta})\cdot P_{t}(\sigma_{2}|S\setminus\{\sigma_{1}\};\bm{\theta})\cdot\;\dots\;\cdot P_{t}(\sigma_{|S|}|\{\sigma_{|S|}\};\bm{\theta})=\prod_{j=1}^{|S|}P_{t}(\sigma_{j}|S^{(j)}_{\sigma};\bm{\theta}),$
Greedy assortment selection for PL loss. Given $H_{t}$ and $\widehat{\bm{\theta}}_{t}$, we define the function $f_{t}(S)$ as follows:
$f_{t}(S):=\frac{1}{|S|}\sum_{j=1}^{|S|}\mathbb{E}_{\begin{subarray}{c}\sigma\sim\mathbb{P}_{t}(\cdot|S;\widehat{\bm{\theta}}_{t})\\ a\sim P_{t}(\cdot|S^{(j)}_{\sigma};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{a^{\prime}\sim P_{t}(\cdot|S^{(j)}_{\sigma};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{H_{t}^{-1}}^{2}\right].$ (10)
Intuitively, finding $S$ that exactly maximizes $f_{t}(S)$ amounts to maximizing the average (mean-centered) uncertainty—that is, the Mahalanobis dispersion of the feature vectors evaluated under $H_{t}^{-1}$. However, computing the exact maximizer of Equation (10) is generally NP-hard. In our setting, it is sufficient to add actions sequentially in a greedy manner. This is because our analysis centers on the suboptimality gap between two actions—one optimal and one chosen by the policy. As a result, the suboptimality gap is (approximately) upper bounded by a term involving $f_{t}(S)$ for pairwise action sets (i.e., $|S|=2$). Therefore, greedily adding actions that increase $f_{t}(S)$ is sufficient.
We initialize $S$ with a pair of actions ($|S|=2$) that maximizes the average information gain, as defined in Equation (10). Then, iteratively add one action at a time by
$a^{\star}\in\mathop{\mathrm{argmax}}_{a\in\mathcal{A}\setminus S}\ \Delta_{t}(a\mid S),\qquad\text{where }\,\Delta_{t}(a\mid S):=f_{t}(S\cup\{a\})-f_{t}(S),$ (11)
[…]
Greedy assortment selection for RB loss. Analogous to Equation (10), we define the following function to select an assortment that maximizes the average uncertainty:
$f_{t}(S)$ $:=\frac{1}{|S|}\sum_{a,a^{\prime}\in S}\mathbb{E}_{\bar{a}\sim P_{t}(\cdot|\{a,a^{\prime}\};\widehat{\bm{\theta}}_{t})}\left[\left\|\phi(x_{t},\bar{a})-\mathbb{E}_{\tilde{a}\sim P_{t}(\cdot|\{a,a^{\prime}\};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},\tilde{a})]\right\|_{H_{t}^{-1}}^{2}\right]$
$=\frac{1}{2|S|}\sum_{a,a^{\prime}\in S}\dot{\mu}\left((\phi(x_{t},a)-\phi(x_{t},a^{\prime}))^{\top}\widehat{\bm{\theta}}_{t}\right)\left\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\right\|_{H_{t}^{-1}}^{2}.$ (12)
[…]
Once the assortment $S_{t}$ is selected, the algorithm receives the ranking feedback $\sigma_{t}$ from the labeler (Line 5) and updates the parameter using Procedure 1 for the PL loss or Procedure 2 for the RB loss (Line 6). After $T$ rounds, the algorithm returns the final policy $\widehat{\pi}_{T}$, which selects actions by maximizing the estimated reward under the final parameter estimate $\widehat{\bm{\theta}}_{T+1}$ (Line 7).
@@ 5.1 Suboptimality Gap of M-AUPO
We begin by presenting the online confidence bound for the PL loss, derived by extending the results of lee2025improved, who analyzed the MNL model [mcfadden1977modelling]. Since the PL model constructs ranking probabilities as a product of MNL probabilities, their confidence bound can be directly applied to our setting by replacing the round $t$ with the cumulative number of updates $\sum_{s=1}^{t}|S_{s}|$.
@@
This confidence bound is free of any polynomial dependency on $K$, which is primarily made possible by the improved self-concordant-like properties proposed by lee2025improved. Moreover, for the RB loss, we can derive a confidence bound of the same order (see Corollary D.1). Based on this confidence bound, we derive the suboptimality gap for M-AUPO, with the proof deferred to Appendix C.
Theorem 1 (Suboptimality gap for PL loss).
Let $\delta\in(0,1]$. Set $\lambda=\Omega\big(d\log(KT/\delta)+\eta(B+d)\big)$ and $\eta=\frac{1}{2}(1+3\sqrt{2}B)$. Define $\kappa:=e^{-6B}$. If Assumption 1 holds, then for any $T\geq e^{K/d}$, with probability at least $1-\delta$, M-AUPO (Algorithm 3) achieves the following suboptimality gap:
$\textup{{SubOpt}}(T)=\tilde{\mathcal{O}}\left(\frac{d}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}+\frac{d^{2}K^{4}}{\kappa T}\right).$
Discussion of Theorem 1. For sufficiently large $T$, the second (non-leading) term becomes negligible, and Theorem 1 shows that the suboptimality gap of M-AUPO decreases as the assortment size $|S_{t}|$ increases. This establishes a strict advantage of receiving ranking feedback over larger assortments. Moreover, our result does not involve any $\mathcal{O}(e^{B})$ dependency in the leading term, a harmful dependency that commonly appears in prior works [saha2021optimal, saha2023dueling, zhu2023principled, xiong2023iterative, zhan2023provable, das2024active, thekumparampil2024comparing, kveton2025active]. Although very recent studies [di2024nearly, chen2025avoiding] also achieve $\mathcal{O}(e^{B})$-free performance in the leading term, they rely on auxiliary techniques and are restricted to pairwise preference feedback. To the best of our knowledge, this is the first theoretical study that simultaneously establishes (i) the performance benefits of utilizing richer ranking feedback over larger assortments, and (ii) the elimination of the $\mathcal{O}(e^{B})$ dependence in the leading term of the PbRL framework when accommodating multiple (i.e., more than two) options.
Proof sketch of Theorem 1. We provide a proof sketch for Theorem 1. For simplicity, we define $\psi_{t,a,a^{\prime}}:=\phi(x_{t},a)-\phi(x_{t},a^{\prime})$. We begin by defining the set of warm-up rounds, denoted by ${\mathcal{T}}^{w}$, which consists of rounds with large (non-centered) uncertainty:
${\mathcal{T}}^{w}$ $:=\left\{t\in[T]:\max_{a,a^{\prime}\in\mathcal{A}}\|\psi_{t,a,a^{\prime}}\|_{H_{t}^{-1}}\geq 1/\left(3\sqrt{2}K\beta_{T+1}(\delta)\right)\right\}.$
@@
$\textup{{SubOpt}}(T)$ $=\frac{1}{T}\sum_{t=1}^{T}\underbrace{\big(\psi_{t,\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})}\big)^{\top}\bm{\theta}^{\star}}_{\text{realized regret of $\widehat{\pi}_{T}$ at round }t}+\underbrace{\frac{1}{T}\sum_{t=1}^{T}\text{MDS}_{t}}_{=\tilde{\mathcal{O}}(1/\sqrt{T})}$
@@
$\lesssim\frac{1}{T}\sum_{t\notin{\mathcal{T}}^{w}}\|\psi_{t,\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})}\|_{H_{t}^{-1}}\underbrace{\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\|_{H_{T}}}_{\leq\beta_{t}(\delta)=\tilde{\mathcal{O}}(\sqrt{d})}+\tilde{\mathcal{O}}\left(\frac{1}{\sqrt{T}}+\frac{d^{2}K^{4}}{\kappa T}\right).$
@@
In the inequality, we first use the fact that $\big(\psi_{t,\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})}\big)^{\top}\widehat{\bm{\theta}}_{T+1}\geq 0$, which follows from definition of $\widehat{\pi}_{T}$. We then apply Hölder’s inequality together with the inequality $H_{T+1}\succeq H_{t}$. Moreover, we invoke Lemma C.4, which states that the regret incurred during the warm-up rounds ${\mathcal{T}}^{w}$ is a non-leading term. Finally, by applying Corollary 1, we bound $\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\|_{H_{T}}$ by $\tilde{\mathcal{O}}(\sqrt{d})$.
2) Avoiding $\mathcal{O}(e^{B})$ for non-warm-up rounds. For $t\notin{\mathcal{T}}^{w}$, we have
$\big(\psi_{t,\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})}\big)^{\top}\bm{\theta}^{\star}\leq\|\psi_{t,\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})}\|_{H_{t}^{-1}}\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\|_{H_{T}}\leq\frac{\beta_{t}(\delta)}{3\sqrt{2}K\beta_{T+1}(\delta)}\leq 1,$
which implies that $1/\dot{\mu}\big(\big(\psi_{t,\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})}\big)^{\top}\bm{\theta}^{\star}\big)\leq(1+e)^{2}$. Thus, we obtain
$\frac{1}{T}\sum_{t\notin{\mathcal{T}}^{w}}\|\psi_{t,\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})}\|_{H_{t}^{-1}}$ $\leq\frac{(1+e)^{2}}{T}\sum_{t\notin{\mathcal{T}}^{w}}\dot{\mu}\big(\big(\psi_{t,\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})}\big)^{\top}\bm{\theta}^{\star}\big)\|\psi_{t,\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})}\|_{H_{t}^{-1}}$
$\leq\frac{e(1+e)^{2}}{T}\sum_{t\notin{\mathcal{T}}^{w}}\dot{\mu}\big(\big(\psi_{t,\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})}\big)^{\top}\widehat{\bm{\theta}}_{t+1}\big)\|\psi_{t,\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})}\|_{H_{t}^{-1}},$
where the last inequality follows from Lemma E.2, together with the fact that for any $a,a^{\prime}\in\mathcal{A}$ and $t\notin{\mathcal{T}}^{w}$, $|\psi_{t,a,a^{\prime}}^{\top}(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{t+1})|\leq\|\psi_{t,a,a^{\prime}}\|_{H_{t}^{-1}}\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{t+1}\|_{H_{T}}\leq\beta_{t}(\delta)/\beta_{T+1}(\delta)=1$. Then, by the Cauchy–Schwarz inequality, and ignoring constants, we have
$\frac{1}{T}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}|S_{t}|\cdot\dot{\mu}\big(\big(\psi_{t,\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})}\big)^{\top}\widehat{\bm{\theta}}_{t+1}\big)\|\psi_{t,\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})}\|_{H_{t}^{-1}}^{2}}$
$\lesssim\frac{1}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}|S_{t}|\cdot f_{t}(\{\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})\})}\leq\frac{1}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}|S_{t}|\cdot f_{t}(S_{t})}$ ($S_{t}$ selection)
$\lesssim\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{t}|}\mathbb{E}_{\begin{subarray}{c}\mathbb{P}_{t}(\cdot|S_{t};\bm{\theta}^{\star})\\ P_{t}(\cdot|S_{t}^{(j)};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{P_{t}(\cdot|S_{t}^{(j)};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{M_{t}^{-1}}^{2}\right]},$ (Lemma C.2)
where $M_{t}=\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{s}|}\mathbb{E}_{\sigma\sim\mathbb{P}_{s}(\cdot|S_{s},\bm{\theta}^{\star})}\left[\nabla^{2}\ell^{(j)}_{s,\sigma}(\widehat{\bm{\theta}}_{s})\right]+\lambda\mathbf{I}_{d}$ is the expected version of $H_{t}$. Finally, applying the elliptical potential lemma (Lemma C.3), we concludes the proof.
Remark 2 (Generality of our technique for avoiding $\mathcal{O}(e^{B})$).
Our approach of dividing the total rounds into warm-up and non–warm-up phases to improve the $\mathcal{O}(e^{B})$ dependency is broadly applicable. In particular, this technique can be readily incorporated into most existing PbRL and dueling bandit algorithms without modifying their original algorithms. See Appendix G.1 for more details.
Furthermore, under the rank-breaking (RB) parameter update (7) and assortment selection rule (12), we establish a comparable suboptimality gap for the RB loss. The proof is provided in Appendix D.
Theorem 2 (Online confidence bound for RB loss).